ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 123, 2001

Abraham, U., M. Rubin and R. Bonnet On a superatomic Boolean algebra which is not generated by a well-founded sublattice

221

Akcoglu, M. and J. Baxter Tail field representations and the zero-two law

253

Amour, L. and T. Raoux L^{1} decay properties for a semilinear parabolic system

157

Baxter, J. See M. Akcoglu

Beidar, K. I. and M. A. Chebotar On Lie derivations of Lie ideals of prime algebras

131

Bonnet, R. See U. Abraham

Carbone, A. Cycles of relatively prime length and the road coloring problem

303

Chebotar, M. A. See K. I. Beidar

Derriennic, Y. and M. Lin Fractional Poisson equations and ergodic theorems for fractional coboundaries

93

Edmunds, D. E. and R. Hurri-Syrjänen Sobolev inequalities of exponential type

61

Einstein-Matthews, S. M. and C. H. Lutterodt Approximation and interpolation on complex algebraic varieties in pseudoconvex domains

189

Gao, F. Metric entropy of convex hulls  359
Ginzburg, D. and D. Jiang Periods and liftings: From G_{2} to C_{3}  29

Hurri-Syrjänen, R. See D. E. Edmunds

Jiang, D. See D. Ginzburg

Kim, J.-L. and I. I. Piatetski-Shapiro Quadratic base change of

θ_{10} 317

Li, B. Q. and E. Villamor Interpolation in the unit ball of C^{n}  341

Lin, M. See Y. Derriennic

Lutterodt, C. H. See S. M. Einstein-Matthews

Maleva, O. A pathological example of a uniform quotient mapping between Euclidean spaces

211

Martínez, J. C. A consistency result on thin-very tall Boolean algebras 273

Matoušek, J. and Y. Rabinovich On dominated l_{1} metrics  285

Mednykh, A. and M. Reni Twofold unbranched coverings of genus two 3-manifolds are hyperelliptic

149

Paoletti, R. Symplectic submanifolds in special position  241

Paternain, G. P. Schrödinger operators with magnetic fields and minimal action functionals

1

Piatetski-Shapiro, I. I. See J.-L. Kim

Rabinovich, Y. See J. Matoušek

Raoux, T. See L. Amour

Reni, M. See A. Mednykh

Riley, D. M. Infinitesimally PI radical algebras  365

Rubin, M. See U. Abraham

Shepherd-Barron, N. I. On simple groups and simple singularities 179

Villamor, E. See B. Q. Li